# A lower bound for the Weisfeiler-Leman dimension of circulant graphs

### v1

@@ 5.1. S-rings
Let $G$ be a finite group. A free ${\mathbb{Z}}$-submodule ${\cal A}$ of the group ring ${\mathbb{Z}}G$ is called a Schur ring (S-ring, for short) over $G$ if there exists a partition ${\cal S}={\cal S}({\cal A})$ of $G$ such that ${\cal A}=\operatorname{Span}_{\mathbb{Z}}\{{\underline{X}}:\ X\in{\cal S}\}$ and
@@
    ${\underline{X}}\,{\underline{Y}}=\sum_{Z\in{\cal S}}c_{X,Y}^{Z}{\underline{Z}}$ for all $X,Y\in{\cal S}$ and some nonnegative integers $c_{X,Y}^{Z}$.
The elements of ${\cal S}$ and the number $\operatorname{rk}({\cal A})=|{\cal S}|$ are called, respectively, the basic sets and rank of ${\cal A}$. The integer $c_{X,Y}^{Z}$ is equal to the number of representations $z=xy$ with $x\in X$ and $y\in Y$ for a fixed $z\in Z$. For any subgroup $M\leq\operatorname{Aut}(G)$, the set $\{x^{M}:\ x\in G\}$ of all $M$-orbits forms a partition ${\cal S}({\cal A})$ for some S-ring ${\cal A}$ over $G$; we denote this S-ring by $\operatorname{cyc}(M,G)$ and call it cyclotomic.
The partial order $\leq$ on the S-rings over $G$ is induced by inclusion. Thus, ${\cal A}\leq{\cal A}^{\prime}$ if and only if any basic set of ${\cal A}$ is a union of some basic sets of ${\cal A}^{\prime}$; in this case, we say that ${\cal A}^{\prime}$ is an extension of ${\cal A}$. The least and greatest elements with respect to $\leq$ are, respectively, the trivial S-ring spanned by ${\underline{1_{G}}}$ and ${\underline{G}}$, and the group ring ${\mathbb{Z}}G$. For any $X\subseteq G$, we denote by $\operatorname{WL}_{G}({\underline{X}})$ the smallest S-ring over $G$, that contains the element ${\underline{X}}$.
@@ 5.2. Isomorphisms
Let ${\cal A}$ and ${\cal A}^{\prime}$ be S-rings over groups $G$ and $G^{\prime}$, respectively, and let $f:G\to G^{\prime}$ be a bijection. We say that $f$ is a (combinatorial) isomorphism from ${\cal A}$ to ${\cal A}^{\prime}$ if for each $X\in{\cal S}({\cal A})$, there is $X^{\prime}\in{\cal S}({\cal A}^{\prime})$ such that $f$ is a graph isomorphism from $\operatorname{Cay}(G,X)$ to $\operatorname{Cay}(G^{\prime},X^{\prime})$, i.e.,
@@
for all $x,y\in G$, where $x^{\prime}=x^{f}$ and $y^{\prime}=y^{f}$. [^3] This condition means exactly that $(Xx)^{f}=X^{\prime}x^{\prime}$ for all $X\in{\cal S}$ and $x\in G$. The isomorphism $f$ is said to be normalized if $f(1_{G})=1_{G^{\prime}}$. Note that the S-rings ${\cal A}$ and ${\cal A}^{\prime}$ are isomorphic if and only if there is a normalized isomorphism from ${\cal A}$ to ${\cal A}^{\prime}$.
The group of all isomorphisms from the S-ring ${\cal A}$ to itself is denoted by $\operatorname{Iso}({\cal A})$. It has a normal subgroup $\operatorname{Aut}({\cal A})$, the automorphism group of ${\cal A}$, that is equal to the intersection of the groups $\operatorname{Aut}(\operatorname{Cay}(G,X))$, $X\in{\cal S}$. In particular,
(10) $G_{right}\leq\operatorname{Aut}({\cal A})\leq\operatorname{Aut}(\operatorname{Cay}(G,X))$
for each $X$. Thus, $f\in\operatorname{Aut}({\cal A})$ if and only if for all $x,y\in G$ and all $X\in{\cal S}$, the elements $y\,x^{-1}$ and $y^{\prime}\,{x^{\prime}}^{-1}$ belong or do not belong to $X$ simultaneously. Note that if $f\in\operatorname{Aut}(G)$ or $f$ is a normalized isomorphism of ${\cal A}$, then $f\in\operatorname{Aut}({\cal A})$ if and only if $f$ leaves every basic set of ${\cal A}$ fixed.
5.3. ${\cal A}$-sets
Any union of basic sets is called an ${\cal A}$-set. The set of all of them is closed with respect to taking inverse, product, and standard set-theoretical operations. An ${\cal A}$-set which is a subgroup of $G$ is called an ${\cal A}$-subgroup. For example, if $X$ is an ${\cal A}$-set, then the groups $\langle{X}\rangle$ and $\operatorname{rad}(X)$ are ${\cal A}$-subgroups.
@@ 5.3. ${\cal A}$-sets
Let ${\cal A}$ be an S-ring over an abelian group of order $n$, and ${\cal S}={\cal S}({\cal A})$. Then
@@
    $X^{(m)}\in{\cal S}$ for any integer $m$ coprime to $n$ and any $X\in{\cal S}$,
@@
    $X^{[p]}\in{\cal S}^{\cup}$ for any prime $p$ dividing $n$ and any $X\in{\cal S}^{\cup}$. [^4] Though Lemma 2.1 in [11] requires $X$ to be the basic set of ${\cal A}^{\prime}$, the proof of the lemma remains true under a weaker assumption that $X$ is an ${\cal A}^{\prime}$-set.
@@ 5.4. Tensor and wreath products
Let ${\cal A}_{1}$ and ${\cal A}_{2}$ be S-rings over groups $G_{1}$ and $G_{2}$, respectively. The tensor product ${\cal A}_{1}\otimes{\cal A}_{2}$ and the wreath product ${\cal A}_{1}\wr{\cal A}_{2}$ are defined to be the S-rings over the direct product $G=G_{1}\times G_{2}$, such that
${\cal S}({\cal A}_{1}\otimes{\cal A}_{2})=\{X_{1}\times X_{2}:X_{1}\in{\cal S}_{1},\ X_{2}\in{\cal S}_{2}\},$
where ${\cal S}_{1}={\cal S}({\cal A}_{1})$ and ${\cal S}_{2}={\cal S}({\cal A}_{2})$, and
${\cal S}({\cal A}_{1}\wr{\cal A}_{2})=\{X_{1}\times\{1_{G_{2}}\}:X_{1}\in{\cal S}_{1}\}\,\cup\,\{G_{1}\times X_{2}:1_{G_{2}}\neq X_{2}\in{\cal S}_{2}\},$
respectively. If $H_{1}\leq G_{1}$ and $H_{2}\leq G_{2}$ are, respectively, ${\cal A}_{1}$- and ${\cal A}_{2}$-subgroups, then $H_{1}H_{2}$ is an $({\cal A}_{1}\otimes{\cal A}_{2})$-subgroup, whereas $H_{1}$ and $G_{1}H_{2}$ are $({\cal A}_{1}\wr{\cal A}_{2})$-subgroups. If, in addition, the group $G_{1}\times G_{2}$ is cyclic, then these are the only $({\cal A}_{1}\otimes{\cal A}_{2})$- and $({\cal A}_{1}\wr{\cal A}_{2})$-subgroups, respectively.
@@ 5.5. Sections
Let $U$ and $L$ be ${\cal A}$-subgroups and also $L$ a normal subgroup of $U$. The group ${\mathfrak{s}}=U/L$ is called an ${\cal A}$-section; the set of all ${\cal A}$-sections is denoted by ${\mathfrak{S}}({\cal A})$. Furthermore, let $\pi:U\to U/L$ be the natural epimorphism. The set
${\cal S}_{\mathfrak{s}}=\{\pi(X):\ X\in{\cal S},\ X\subseteq U\}$
forms a partition of the group ${\mathfrak{s}}$, and the elements ${\underline{\pi(X)}}$ span an S-ring over ${\mathfrak{s}}$; it is denoted by ${\cal A}_{\mathfrak{s}}$. If $U=\langle{X}\rangle$ and $L=\operatorname{rad}(X)$ for some $X\in{\cal S}$, then the section ${\mathfrak{s}}$ is said to be principal and is denoted by ${\mathfrak{s}}_{X}$. In this case, setting ${\mathfrak{s}}={\mathfrak{s}}_{X}$, $\pi_{\mathfrak{s}}=\pi$, and $X_{\mathfrak{s}}=\pi(X)$, we have
@@
Let ${\mathfrak{s}}=U/L$ and ${\mathfrak{t}}=U^{\prime}/L^{\prime}$ be ${\cal A}$-sections. We say that ${\mathfrak{s}}$ is a subsection of ${\mathfrak{t}}$ if $U\leq U^{\prime}$ and $L\geq L^{\prime}$; in this case, we write ${\mathfrak{s}}\preceq{\mathfrak{t}}$. This defines a partial order on the set ${\mathfrak{S}}({\cal A})$. The greatest element in it is $G/1$; every minimal element is of the form $H/H$, where $H$ is an ${\cal A}$-subgroup.
@@ 5.6. Projective equivalence
An ${\cal A}$-section ${\mathfrak{t}}=U^{\prime}/L^{\prime}$ is called a multiple of the section ${\mathfrak{s}}=U/L$ if
@@
The projective equivalence relation on the set ${\mathfrak{S}}({\cal A})$ of all ${\cal A}$-sections is defined to be the symmetric transitive closure of the relation “to be a multiple”; we write ${\mathfrak{s}}\sim{\mathfrak{t}}$ if the sections ${\mathfrak{s}}$ and ${\mathfrak{t}}$ are projectively equivalent. Any two projectively equivalent sections are obviously isomorphic as groups. Moreover, if ${\mathfrak{s}}\sim{\mathfrak{t}}$, then there is a group isomorphism $f=f_{{\mathfrak{s}},{\mathfrak{t}}}$ from ${\mathfrak{s}}$ to ${\mathfrak{t}}$ belonging to $\operatorname{Iso}({\cal A}_{\mathfrak{s}},{\cal A}_{\mathfrak{t}})$ and such that $(\gamma^{\mathfrak{s}})^{f}=\gamma^{\mathfrak{t}}$ for all normalized automorphisms $\gamma\in\operatorname{Aut}({\cal A})$, where $\gamma^{\mathfrak{s}}$ and $\gamma^{\mathfrak{t}}$ are the restriction of $\gamma$ to ${\mathfrak{s}}$ and to ${\mathfrak{t}}$, respectively (see Subsection 2.1).
@@ 5.7. Algebraic isomorphisms
Let ${\cal A}$ and ${\cal A}^{\prime}$ be S-rings over groups $G$ and $G^{\prime}$, respectively. A ring isomorphism $\varphi:{\cal A}\to{\cal A}^{\prime}$ is called an algebraic isomorphism from ${\cal A}$ to ${\cal A}^{\prime}$, if for any $X\in{\cal S}({\cal A})$, there exists $X^{\prime}\in{\cal S}({\cal A}^{\prime})$ such that
@@
Each normalized isomorphism $f:{\cal A}\to{\cal A}^{\prime}$ induces an algebraic isomorphism for which $X^{\prime}=X^{f}$. However not every algebraic isomorphism is induced by an isomorphism.
From the definition, it follows that the mapping $X\mapsto X^{\prime}$ is a bijection from ${\cal S}({\cal A})$ onto ${\cal S}({\cal A}^{\prime})$. This bijection is naturally extended to a bijection between the ${\cal A}$- and ${\cal A}^{\prime}$-sets, with respect to which the ${\cal A}$-subgroups go to ${\cal A}^{\prime}$-subgroups. Moreover, if the groups $G$ and $G^{\prime}$ are abelian, then the last bijection can be extended to the bijection between the ${\cal A}$- and ${\cal A}^{\prime}$-sections. The images of an ${\cal A}$-set $X$ and an ${\cal A}$-section ${\mathfrak{s}}$ under the corresponding bijections are denoted by $X^{\varphi}$ and ${\mathfrak{s}}^{\varphi}$, respectively. For any such ${\mathfrak{s}}$, the algebraic isomorphism $\varphi$ induces an algebraic isomorphism
$\varphi_{\mathfrak{s}}:{\cal A}_{{\mathfrak{s}}}\to{\cal A}^{\prime}_{{\mathfrak{s}}^{\prime}},\ \pi_{{\mathfrak{s}}}(X)\mapsto\pi_{{\mathfrak{s}}^{\prime}}(X^{\prime}),$
@@
The above bijection between the ${\cal A}$- and ${\cal A}^{\prime}$-sets is in fact an isomorphism of the corresponding lattices. Furthermore, given an ${\cal A}$-set $X$, we have
@@
Let $f:G\to G^{\prime}$ be a bijection taking the identity of $G$ to that of $G^{\prime}$. Then $f$ is a normalized isomorphism from ${\cal A}$ to ${\cal A}^{\prime}$ if and only if there exists an algebraic isomorphism $\varphi:{\cal A}\to{\cal A}^{\prime}$ such that $X^{\varphi}=X^{f}$ for all $X\in{\cal S}({\cal A})$. In this case, we say that $f$ induces $\varphi$. The set of all normalized isomorphisms that induce a fixed $\varphi$ is denoted by $\operatorname{Iso}({\cal A},{\cal A}^{\prime},\varphi)$; we abbreviate $\operatorname{Iso}({\cal A},\varphi)=\operatorname{Iso}({\cal A},{\cal A},\varphi)$.

### v2

@@ 5.1. S-rings
Let $G$ be a finite group. A free $\mathbb{Z}$}-submodule $\mathcal{A}$} of the group ring ${\mathbb{Z}}G$ is called a Schur ring (S-ring, for short) over $G$ if there exists a partition ${\mathcal{S}}={\mathcal{S}}({\mathcal{A}})$ of $G$ such that ${\mathcal{A}}=\operatorname{Span}_{\mathbb{Z}}\{{\underline{X}}:\ X\in{\mathcal{S}}\}$ and
@@
    ${\underline{X}}\,{\underline{Y}}=\sum_{Z\in{\mathcal{S}}}c_{X,Y}^{Z}{\underline{Z}}$ for all $X,Y\in{\mathcal{S}}$ and some nonnegative integers $c_{X,Y}^{Z}$.
The elements of $\mathcal{S}$} and the number $\operatorname{rk}({\mathcal{A}})=|{\mathcal{S}}|$ are called, respectively, the basic sets and rank of $\mathcal{A}$}. The integer $c_{X,Y}^{Z}$ is equal to the number of representations $z=xy$ with $x\in X$ and $y\in Y$ for a fixed $z\in Z$. For any subgroup $M\leq\operatorname{Aut}(G)$, the set $\{x^{M}:\ x\in G\}$ of all $M$-orbits forms a partition ${\mathcal{S}}({\mathcal{A}})$ for some S-ring $\mathcal{A}$} over $G$; we denote this S-ring by $\operatorname{cyc}(M,G)$ and call it cyclotomic.
The partial order $\leq$ on the S-rings over $G$ is induced by inclusion. Thus, ${\mathcal{A}}\leq{\mathcal{A}}^{\prime}$ if and only if any basic set of $\mathcal{A}$} is a union of some basic sets of $\mathcal{A}$$^{\prime}$; in this case, we say that $\mathcal{A}$$^{\prime}$ is an extension of $\mathcal{A}$}. The least and greatest elements with respect to $\leq$ are, respectively, the trivial S-ring spanned by $\underline{1_{G}}$} and $\underline{G}$}, and the group ring ${\mathbb{Z}}G$. For any $X\subseteq G$, we denote by $\operatorname{WL}_{G}({\underline{X}})$ the smallest S-ring over $G$, that contains the element $\underline{X}$}.
@@ 5.2. Isomorphisms
Let $\mathcal{A}$} and $\mathcal{A}$$^{\prime}$ be S-rings over groups $G$ and $G^{\prime}$, respectively, and let $f:G\to G^{\prime}$ be a bijection. We say that $f$ is a (combinatorial) isomorphism from $\mathcal{A}$} to $\mathcal{A}$$^{\prime}$ if for each $X\in{\mathcal{S}}({\mathcal{A}})$, there is $X^{\prime}\in{\mathcal{S}}({\mathcal{A}}^{\prime})$ such that $f$ is a graph isomorphism from $\operatorname{Cay}(G,X)$ to $\operatorname{Cay}(G^{\prime},X^{\prime})$, i.e.,
@@
for all $x,y\in G$, where $x^{\prime}=x^{f}$ and $y^{\prime}=y^{f}$. [^3] This condition means exactly that $(Xx)^{f}=X^{\prime}x^{\prime}$ for all $X\in{\mathcal{S}}$ and $x\in G$. The isomorphism $f$ is said to be normalized if $f(1_{G})=1_{G^{\prime}}$. Note that the S-rings $\mathcal{A}$} and $\mathcal{A}$$^{\prime}$ are isomorphic if and only if there is a normalized isomorphism from $\mathcal{A}$} to $\mathcal{A}$$^{\prime}$.
The group of all isomorphisms from the S-ring $\mathcal{A}$} to itself is denoted by $\operatorname{Iso}({\mathcal{A}})$. It has a normal subgroup $\operatorname{Aut}({\mathcal{A}})$, the automorphism group of $\mathcal{A}$}, that is equal to the intersection of the groups $\operatorname{Aut}(\operatorname{Cay}(G,X))$, $X\in{\mathcal{S}}$. In particular,
(10) $G_{right}\leq\operatorname{Aut}({\mathcal{A}})\leq\operatorname{Aut}(\operatorname{Cay}(G,X))$
for each $X$. Thus, $f\in\operatorname{Aut}({\mathcal{A}})$ if and only if for all $x,y\in G$ and all $X\in{\mathcal{S}}$, the elements $y\,x^{-1}$ and $y^{\prime}\,{x^{\prime}}^{-1}$ belong or do not belong to $X$ simultaneously. Note that if $f\in\operatorname{Aut}(G)$ or $f$ is a normalized isomorphism of $\mathcal{A}$}, then $f\in\operatorname{Aut}({\mathcal{A}})$ if and only if $f$ leaves every basic set of $\mathcal{A}$} fixed.
5.3. $\mathcal{A}$}-sets
Any union of basic sets is called an $\mathcal{A}$}-set. The set of all of them is closed with respect to taking inverse, product, and standard set-theoretical operations. An $\mathcal{A}$}-set which is a subgroup of $G$ is called an $\mathcal{A}$}-subgroup. For example, if $X$ is an $\mathcal{A}$}-set, then the groups $\langle{X}\rangle$ and $\operatorname{rad}(X)$ are $\mathcal{A}$}-subgroups.
@@ 5.3. $\mathcal{A}$}-sets
Let $\mathcal{A}$} be an S-ring over an abelian group of order $n$, and ${\mathcal{S}}={\mathcal{S}}({\mathcal{A}})$. Then
@@
    $X^{(m)}\in{\mathcal{S}}$ for any integer $m$ coprime to $n$ and any $X\in{\mathcal{S}}$,
@@
[…]
@@ 5.4. Tensor and wreath products
Let ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ be S-rings over groups $G_{1}$ and $G_{2}$, respectively. The tensor product ${\mathcal{A}}_{1}\otimes{\mathcal{A}}_{2}$ and the wreath product ${\mathcal{A}}_{1}\wr{\mathcal{A}}_{2}$ are defined to be the S-rings over the direct product $G=G_{1}\times G_{2}$, such that
${\mathcal{S}}({\mathcal{A}}_{1}\otimes{\mathcal{A}}_{2})=\{X_{1}\times X_{2}:X_{1}\in{\mathcal{S}}_{1},\ X_{2}\in{\mathcal{S}}_{2}\},$
where ${\mathcal{S}}_{1}={\mathcal{S}}({\mathcal{A}}_{1})$ and ${\mathcal{S}}_{2}={\mathcal{S}}({\mathcal{A}}_{2})$, and
${\mathcal{S}}({\mathcal{A}}_{1}\wr{\mathcal{A}}_{2})=\{X_{1}\times\{1_{G_{2}}\}:X_{1}\in{\mathcal{S}}_{1}\}\,\cup\,\{G_{1}\times X_{2}:1_{G_{2}}\neq X_{2}\in{\mathcal{S}}_{2}\},$
respectively. If $H_{1}\leq G_{1}$ and $H_{2}\leq G_{2}$ are, respectively, ${\mathcal{A}}_{1}$- and ${\mathcal{A}}_{2}$-subgroups, then $H_{1}H_{2}$ is an $({\mathcal{A}}_{1}\otimes{\mathcal{A}}_{2})$-subgroup, whereas $H_{1}$ and $G_{1}H_{2}$ are $({\mathcal{A}}_{1}\wr{\mathcal{A}}_{2})$-subgroups. If, in addition, the group $G_{1}\times G_{2}$ is cyclic, then these are the only $({\mathcal{A}}_{1}\otimes{\mathcal{A}}_{2})$- and $({\mathcal{A}}_{1}\wr{\mathcal{A}}_{2})$-subgroups, respectively.
@@ 5.5. Sections
Let $U$ and $L$ be $\mathcal{A}$}-subgroups and also $L$ a normal subgroup of $U$. The group ${\mathfrak{s}}=U/L$ is called an $\mathcal{A}$}-section; the set of all $\mathcal{A}$}-sections is denoted by ${\mathfrak{S}}({\mathcal{A}})$. Furthermore, let $\pi:U\to U/L$ be the natural epimorphism. The set
${\mathcal{S}}_{\mathfrak{s}}=\{\pi(X):\ X\in{\mathcal{S}},\ X\subseteq U\}$
forms a partition of the group $\mathfrak{s}$}, and the elements $\underline{\pi(X)}$} span an S-ring over $\mathfrak{s}$}; it is denoted by ${\mathcal{A}}_{\mathfrak{s}}$. If $U=\langle{X}\rangle$ and $L=\operatorname{rad}(X)$ for some $X\in{\mathcal{S}}$, then the section $\mathfrak{s}$} is said to be principal and is denoted by ${\mathfrak{s}}_{X}$. In this case, setting ${\mathfrak{s}}={\mathfrak{s}}_{X}$, $\pi_{\mathfrak{s}}=\pi$, and $X_{\mathfrak{s}}=\pi(X)$, we have
@@
Let ${\mathfrak{s}}=U/L$ and ${\mathfrak{t}}=U^{\prime}/L^{\prime}$ be $\mathcal{A}$}-sections. We say that $\mathfrak{s}$} is a subsection of $\mathfrak{t}$} if $U\leq U^{\prime}$ and $L\geq L^{\prime}$; in this case, we write ${\mathfrak{s}}\preceq{\mathfrak{t}}$. This defines a partial order on the set ${\mathfrak{S}}({\mathcal{A}})$. The greatest element in it is $G/1$; every minimal element is of the form $H/H$, where $H$ is an $\mathcal{A}$}-subgroup.
@@ 5.6. Projective equivalence
An $\mathcal{A}$}-section ${\mathfrak{t}}=U^{\prime}/L^{\prime}$ is called a multiple of the section ${\mathfrak{s}}=U/L$ if
@@
The projective equivalence relation on the set ${\mathfrak{S}}({\mathcal{A}})$ of all $\mathcal{A}$}-sections is defined to be the symmetric transitive closure of the relation “to be a multiple”; we write ${\mathfrak{s}}\sim{\mathfrak{t}}$ if the sections $\mathfrak{s}$} and $\mathfrak{t}$} are projectively equivalent. Any two projectively equivalent sections are obviously isomorphic as groups. Moreover, if ${\mathfrak{s}}\sim{\mathfrak{t}}$, then there is a group isomorphism $f=f_{{\mathfrak{s}},{\mathfrak{t}}}$ from $\mathfrak{s}$} to $\mathfrak{t}$} belonging to $\operatorname{Iso}({\mathcal{A}}_{\mathfrak{s}},{\mathcal{A}}_{\mathfrak{t}})$ and such that $(\gamma^{\mathfrak{s}})^{f}=\gamma^{\mathfrak{t}}$ for all normalized automorphisms $\gamma\in\operatorname{Aut}({\mathcal{A}})$, where $\gamma^{\mathfrak{s}}$ and $\gamma^{\mathfrak{t}}$ are the restriction of $\gamma$ to $\mathfrak{s}$} and to $\mathfrak{t}$}, respectively (see Subsection 2.1).
@@ 5.7. Algebraic isomorphisms
Let $\mathcal{A}$} and $\mathcal{A}$$^{\prime}$ be S-rings over groups $G$ and $G^{\prime}$, respectively. A ring isomorphism $\varphi:{\mathcal{A}}\to{\mathcal{A}}^{\prime}$ is called an algebraic isomorphism from $\mathcal{A}$} to $\mathcal{A}$$^{\prime}$, if for any $X\in{\mathcal{S}}({\mathcal{A}})$, there exists $X^{\prime}\in{\mathcal{S}}({\mathcal{A}}^{\prime})$ such that
@@
Each normalized isomorphism $f:{\mathcal{A}}\to{\mathcal{A}}^{\prime}$ induces an algebraic isomorphism for which $X^{\prime}=X^{f}$. However not every algebraic isomorphism is induced by an isomorphism.
[…]
$\varphi_{\mathfrak{s}}:{\mathcal{A}}_{{\mathfrak{s}}}\to{\mathcal{A}}^{\prime}_{{\mathfrak{s}}^{\prime}},\ \pi_{{\mathfrak{s}}}(X)\mapsto\pi_{{\mathfrak{s}}^{\prime}}(X^{\prime}),$
@@
[…]
@@
Let $f:G\to G^{\prime}$ be a bijection taking the identity of $G$ to that of $G^{\prime}$. Then $f$ is a normalized isomorphism from $\mathcal{A}$} to $\mathcal{A}$$^{\prime}$ if and only if there exists an algebraic isomorphism $\varphi:{\mathcal{A}}\to{\mathcal{A}}^{\prime}$ such that $X^{\varphi}=X^{f}$ for all $X\in{\mathcal{S}}({\mathcal{A}})$. In this case, we say that $f$ induces $\varphi$. The set of all normalized isomorphisms that induce a fixed $\varphi$ is denoted by $\operatorname{Iso}({\mathcal{A}},{\mathcal{A}}^{\prime},\varphi)$; we abbreviate $\operatorname{Iso}({\mathcal{A}},\varphi)=\operatorname{Iso}({\mathcal{A}},{\mathcal{A}},\varphi)$.